OpenStudy (anonymous):

How many potassium atoms are in 250.0 grams of potassium? 1.66 x 10−24 atoms K 3.85 x 1024 atoms K 1.51 x 1026 atoms K 5.88 x 1027 atoms K

OpenStudy (xishem):

\[250.0g\ K*(\frac{1mol\ K}{39.0983g\ K})*(\frac{6.022 \times {10^{23}}atoms\ K}{1mol\ K})=3.85 \times {10^{24}}atoms\ K\]